OpenStudy (anonymous):

simplify the expression 0.11y cubed + 0.6y cubed - .05y cubed.

OpenStudy (anonymous):

just simplify the question there all like terms

OpenStudy (anonymous):

I came up with 0.092y

OpenStudy (anonymous):

and sorry the last one was suppose to be 0.5y cubed

OpenStudy (anonymous):

i got 1.31y

OpenStudy (anonymous):

\[0.1y{3}+0.6y ^{3}-0.5y ^{3}\]

OpenStudy (anonymous):

0.11+0.6-0.05 = 0.71-0.05 = 0.66 Hence , 0.66y cube

OpenStudy (anonymous):

i tried

OpenStudy (anonymous):

0.1y+0.6y=0.7y 0.7y-0.5y=0.2y so would it be \[0.2y^{3}\]

OpenStudy (anonymous):

actually yea it would

OpenStudy (anonymous):

ok thank you and sorry I had the numbers all mixed up

OpenStudy (anonymous):

np